Circular No. 5777

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 J. Y. Hu, Z. W. Li, X. J. Jiang, and L. F. Wang, Beijing Observatory,

e-mail: “Low-resolution spectra (0.53 nm/pixel) of SN 1993J obtained at the Beijing Astronomical Observatory’s 2.16-m telescope show conspicuous changes in the P-Cyg profiles of Hα and Na I 589.3-nm. From Apr. 22 to 26, the shapes of the bottoms of the absorption troughs have changed to sharp triangle-like features from their original flat bottom structure. Higher-resolution spectra (0.13 nm/pixel) show that the absorption trough of the Na 589.3-nm line consists of two components that differ in velocity by ∼ 3000 km/s, of which the bluer one has not changed very much, while the redder component has been evolving rapidly since Apr. 22 and is much stronger than the bluer component as of Apr. 26. The Hα line shows very similar changes. The depletion of the Na 589.3-nm line is ∼ 50 percent, and that of Hα is ∼ 60 percent. These changes in the line profiles may be due to a change in the optical properties of the atmosphere surrounding the supernova’s photosphere, suggesting perhaps an asymmetric explosion.”

Z.-Y. Zheng, X. Zhou, and H. Wu, Beijing Astronomical Observatory, provide the following data as a total replacement for their erroneous photometry published on IAUC 5774 (the revised reduction was performed using the data from Corwin on IAUC 5742; internal errors ∼ ± 0.02): Apr. 7.56 UT, U = 11.81, B = 12.15, V = 11.66, R = 11.30, I = 11.14; 9.61, 11.77, 12.05, 11.39, 11.07, 10.92; 10.58, 11.69, 11.82, 11.27, 10.97, 10.84; 11.59, 11.54, 11.75, 11.17, 10.81, 10.74; 12.64, 11.47, 11.68, 11.06, 10.73, 10.66; 13.61, 11.39, 11.56, 11.02, 10.69, 10.59; 14.62, 11.37, 11.51, 10.92, 10.57, 10.49; 15.59, 11.39, 11.48, 10.86, 10.55, 10.46; 16.55, 11.35, 11.49, 10.79, 10.50, 10.42; 17.55, 11.39, 11.42, 10.84, 10.48, 10.39; 18.57, 11.43, 11.51, 10.81, 10.48, 10.39; 19.58, 11.56, 11.55, 10.82, 10.46, 10.38; 20.55, 11.69, 11.67, 10.86, 10.49, 10.35; 21.56, 11.85, 11.83, 10.91, 10.52, 10.40; 23.56, 12.29, 12.07, 11.06, 10.60, 10.48; 24.56, 12.57, 12.29, 11.15, 10.66, 10.53; 25.56, 12.81, 12.51, 11.28, 10.76, 10.59; 26.54, 13.04, 12.63, 11.30, 10.79, 10.60; 29.54, 13.60, 12.90, 11.57, 10.99, 10.80.

NOVA OPHIUCHI 1993 Visual magnitude estimates: Apr. 25.81 UT, 9.8 (P. Camilleri, Cobram,

Vic., Australia); 26.45, 10.8 (R. Royer, Lakewood, CA); 27.05, 11.6 (A. Mizser, Budapest, Hungary); 29.40, 11.9 (Royer).

1993 April 30 Daniel W. E. Green